what should the radius of a cone be with a volume of 8 and a height of 6 be?
Answer by rfer(16322)

V=pir^2h/3
8=(3.14*r^2*6)/3
8*3=18.84r^2
24/18.84=r^2
1.27=r^2
r=1.129 units